# STF: Shallow-Level Temporal Feedback to Enhance Spiking Transformers

### v1

@@ 5 Conclusion
In this paper, we propose Shallow-level Temporal Feedback (STF), a lightweight module that can be seamlessly integrated into the encoding layer of SNNs. Unlike deep-level feedback loops, STF avoids high-dimensional transformations and redundant forward passes. Experimental results show that STF consistently improves the performance of various Transformer-based SNN backbones under different timestep settings, with only a modest increase in the model size, energy consumption, and inference latency. Further analysis reveals that the performance gain can be primarily attributed to the enhanced diversity of the spike patterns. Moreover, we compare STF with direct coding and its variants in terms of adversarial robustness and temporal information, showing that STF is promising as a novel spike encoding scheme for static scenarios. However, additional experiments reveal that STF exhibits suboptimal performance on the neuromorphic dataset. In future work, we will aim to improve its adaptability to event-based scenarios.
@@ Appendix A Dataset Description
CIFAR-10: As a classic benchmark in image classification, CIFAR-10 comprises 60,000 colored images at a resolution of $\times 32$ 32 pixels. The dataset is evenly split into 10 distinct categories, with 6,000 images per class.
CIFAR-100: Sharing the same structure as CIFAR-10, CIFAR-100 presents a more complex challenge by increasing the number of classes to 100. It also consists of 60,000 color images of $\times 32$ 32 pixels, with 600 samples per class, making it suitable for evaluating fine-grained recognition capabilities.
@@ B.1 Spikformer
[…]
@@ B.2 SDT
Table 1: Training hyperparameters for SDT on CIFAR datasets and ImageNet-1K
@@ B.3 QKFormer
[…]
Image size 32 224
[…]
Timestep 2/4/6/8 4
Batch size 64   512$\ddagger$
@@ Appendix D Ablation Study on Variants of the STF Module
Table 3 presents a detailed performance comparison of different STF variants on the CIFAR datasets, corresponding to the visualizations in Figure 2 of the main text. Experiments demonstrate that STF-4 consistently outperforms other STF variants under different time-step settings across various Transformer-based SNN backbones. Unless explicitly specified, STF-4 serves as the default configuration throughout this paper.
@@ Appendix G $SG_{t}$ Derivation Under Static Datasets
From Equation (12) and the geometric series properties, we can obtain:
@@
Table 5: Comparison of ablation studies on TSPE and TF components.
@@ Appendix H Definition of Spike Entropy
where $p_{i}$ denotes the proportion of each distinct spike pattern. Assuming the spike timestep is $T$, the index $i$ ranges from 111 to $2^{T}$, covering all possible patterns. Higher spike entropy reflects greater diversity in neural spike patterns, implying a higher potential information capacity.
@@ Appendix I Adversarial Evaluation and Statistical Results
As shown in Tables 7, 8, and 9, the detailed and comprehensive experimental results demonstrate that STF achieves better robustness without sacrificing clean performance, compared to direct coding.
@@
We conduct detailed ablation experiments on the TSPE and TF components using Spikformer as the backbone. The experiments are performed under four spike-timestep settings on the CIFAR datasets to evaluate their impact on model performance and the diversity of the spike patterns. Following the methodology described in Section H, we use spike entropy as the metric to quantify spike pattern diversity. As shown in Table 5, TF leads to more substantial improvements in both spike pattern diversity and model performance compared to TSPE, while their combination yields further gains.
@@
As shown in Tables 10–19, we compare the spike pattern statistics at the output of LIF neurons in the encoding layer of Spikformer on the CIFAR-10 dataset. The comparison covers direct coding, its variants (GAC coding and IMP coding), and our proposed STF under different time steps ($T=2,4,6,8$). Additional detailed statistics for other backbone architectures can be found in the submitted supplementary material.
See pages - of aaai26_appendix.pdf

### v2

@@ 5 Conclusion
In this paper, we propose Shallow-level Temporal Feedback (STF), a lightweight module that can be seamlessly integrated into the encoding layer of SNNs. Unlike deep-level feedback loops, STF avoids high-dimensional transformations and redundant forward passes. Experimental results show that STF consistently improves the performance of various Transformer-based SNN backbones under different timestep settings, with only a modest increase in the model size, energy consumption, and inference latency. Further analysis reveals that the performance gain can be primarily attributed to the enhanced diversity of spike patterns. Moreover, we compare STF with direct coding and its variants in terms of adversarial robustness and temporal information, showing that STF is promising as a novel spike encoding scheme for static scenarios. However, additional experiments reveal that STF exhibits suboptimal performance on the neuromorphic dataset. In future work, we will aim to improve its adaptability to event-based scenarios.
@@ Appendix A Dataset Description
CIFAR-10: As a classic benchmark in image classification, CIFAR-10 comprises 60,000 colored images at a resolution of $32\times 32$ pixels. The dataset is evenly split into 10 distinct categories, with 6,000 images per class.
CIFAR-100: Sharing the same structure as CIFAR-10, CIFAR-100 presents a more complex challenge by increasing the number of classes to 100. It also consists of 60,000 color images of $32\times 32$ pixels, with 600 samples per class, making it suitable for evaluating fine-grained recognition capabilities.
@@ B.1 Spikformer
We follow the protocol of Zhou et al. (Zhou et al. 2023) and train all models from scratch on the CIFAR-10 and CIFAR-100 datasets for 300 epochs using the AdamW optimizer, with an initial learning rate of 5e45e-4, a weight decay of 6e26e-2, and a batch size of 128. A cosine annealing learning rate scheduler is employed, with 20 warm-up epochs and 10 cool-down epochs. To ensure reproducibility, the random seed is set to 42.
@@ B.2 SDT
Table 5: Training hyperparameters for SDT on the CIFAR datasets and ImageNet-1K
@@ Appendix D Ablation Study on Variants of the STF Module
Table 7 presents a detailed performance comparison of different STF variants on the CIFAR datasets, corresponding to the visualizations in Figure 3 of the main text. Experiments demonstrate that STF-4 consistently outperforms other STF variants under different timestep settings across various Transformer-based SNN backbones. Unless explicitly specified, STF-4 serves as the default configuration throughout this paper.
@@ Appendix G $SG_{t}$ Derivation Under Static Datasets
From Equation 10 and the geometric series properties, we can obtain:
@@
Table 6: Comparison of ablation studies on the TSPE and TF components.
@@ Appendix H Definition of Spike Entropy
where $p_{i}$ denotes the proportion of each distinct spike pattern. Assuming the spike timestep is $T$, the index $i$ ranges from $1$ to $2^{T}$, covering all possible patterns. Higher spike entropy reflects greater diversity in neural spike patterns, implying a higher potential information capacity.
@@ Appendix I Adversarial Evaluation and Statistical Results
As shown in Tables LABEL:tab:spikformer_backbone_adv_appendix, LABEL:tab:spikedriven_backbone_adv_appendix, and LABEL:tab:qkformer_backbone_adv_appendix, the detailed and comprehensive experimental results demonstrate that STF achieves better robustness without sacrificing clean performance, compared to direct coding.
@@
We conduct detailed ablation experiments on the TSPE and TF components using Spikformer as the backbone. The experiments are performed under four timestep settings on the CIFAR datasets to evaluate their impact on model performance and the diversity of spike patterns. Following the methodology described in Section H, we use spike entropy as the metric to quantify the diversity of spike patterns. As shown in Table 6, TF leads to more substantial improvements in both the diversity of spike patterns and model performance compared to TSPE, while their combination yields further gains.
@@
As shown in Tables LABEL:tab:spikformer_spike_patterns_all_cifar10_t_2LABEL:tab:spikformer_spike_pattern_cifar10_STF_coding_t_8, we compare the spike pattern statistics at the output of LIF neurons in the encoding layer of Spikformer on the CIFAR-10 dataset. The comparison covers direct coding, its variants (GAC coding and IMP coding), and our proposed STF under different time steps ($T=2,4,6,8$). Additional detailed statistics for other backbone architectures can be found in the submitted supplementary material.
See pages 1-19 of aaai26_appendix.pdf